# Dredge Word, Social Media, and Webgraph Networks for Unreliable Website Classification and Identification

### v1

@@ Acknowledgements
7 Limitations
While we attempted to mitigate limitations where possible, several substantial ones exist. First, the Twitter data were collected from 2020-2022, but the dredge words used in this paper were based on Google SERPs in 2024. We cannot say that the dredge words we extracted surfaced the same content in 2024 as when the phrases were initially used on Twitter. However, given the conspiratorial, niche, and pseudo-scientific nature of many of the dredge words we identified, we speculate that they would be unlikely to surface substantial reliable content, particularly as reliable sites have had several years to attempt to debunk—and thereby rank for—dredge word queries. Future work could explore the impacts of better temporal alignment, which would have the added benefit of covering data voids formed around breaking news stories.
The Partial F1 metric we use for discovery evaluation is limited, as previous work has shown many PoliticalNews and MBFC sites on which the metric relies are dead or no longer active Carragher et al. (2024). However, to our knowledge, no better metric exists for evaluating unreliable domain discovery systems. This is a needed avenue for future work. Additionally, the current process for extracting dredge words relies on paid third-party services. While we validated each of the extracted dredge words to ensure they are ranking in Google’s top-10, approaches to extract keyphrases from, for example, common-crawl webgraphs would provide more democratization of this line of research. We publish all dredge word data collected for this project in line with that goal.
8 Ethical Considerations
To avoid unintentionally boosting the search rankings of unreliable domains, we do not directly link to any referenced unreliable domains. For unreliable sites, we link to either archive.org or third-party reliability assessments of those domains. We further acknowledge the inherent challenges in determining the “reliability” of websites, and note that even reliable websites can occasionally publish inaccurate content. We defaulted entirely to third parties in our original acquisition of labels, but emphasize that any reliability-labeling ontology should be scrutinized to minimize harms. We would argue that the system we proposed could be generalized to any reasonable reliability ontology.
Acknowledging that the Twitter data likely contains examples of “sensitive events” which could compromise the privacy or safety of users, we adhere to Twitter’s terms of service and do not release this data [^8] https://developer.x.com/en/developer-terms/agreement-and-policy. Finally, we acknowledge that by publicly releasing these dredge-word lists, bad actors could use these terms to direct people to unreliable information. However, once data voids are identified, they are often plugged Norocel and Lewandowski (2023). We hope that by releasing this data, reliable websites can target these phrases to reduce both the visibility and the ease of access of unreliable content.
Appendix A Dredge Word EDA
A.1 Dredge Word SERPs
To understand the distributions of websites surfaced in SERPs, we will provide an illustrative example. One could imagine a user browsing a social media platform and coming across a tweet or post like the one in Figure 3.
Figure 3: A tweet about “indigo children”.
A user encountering new, potentially perplexing information, might then be curious and compelled to search “indigo children meaning” or just “indigo children” on Google—both phrases are dredge words. We provide the results for former query in Figure 4. We note that Google’s definition snippet expresses pseudo-scientific concepts. YouTube videos, followed by a Reddit post from the /r/Psychic subreddit were the fourth-to-the-seventh highest-ranked URLs when we manually searched this dredge word phrase in June of 2024. Another user who may have searched only “indigo children”, would have seen a snippet from Wikipedia which correctly calls indigo children pseudoscientific. However, the user would have seen four books for sale about indigo children alongside the snippet, and the second highest-ranking return was an article from an unreliable domain identifying the “13 signs you’re an indigo child”. Were the user to then scroll down to the 6th result, they would see an Amazon link to a book on indigo children.
Figure 4: SERP results for “Indigo Children Meaning”
A.2 Dredge EDA
@@
We applied the python library langdetect Danilák (2014) to the list of 3,933 dredge words used in the paper, but we observe that automated language extraction of dredge words is a challenge. We observed many dredge words are disembodied fragments, names of people, or simply foreign words adopted by various communities. For example, several Hindi-language-origin yoga poses appear in the dredge word list: parivrtta anjaneyasana and parivrtta utkatasana. These terms surface gaia, a conspiracy pseudoscience website [^9] https://mediabiasfactcheck.com/gaia/. Some websites rank highly for the same terms in multiple langauges; the arabic and chinese words for (coral) both surface creation, a website that frequently promotes pseudoscience [^10] https://mediabiasfactcheck.com/christian-ministries-international/. Qualitatively, hatespeech is fairly uncommon in the dredge words we identified; 11 unique dredge word phrases contain the non-euphemistic writing of the n-word and 3 others contain a homophobic slur.
Appendix B Curriculum EDA
@@
Among the websites rated most reliable, the style, content, and widespread name recognition would very likely result in most annotators to correctly identifying the sites as reliable (e.g., reuters.com, nasa.gov, smithsonianmag.com, nature.com). The landing pages of some of the least reliable labeled domains are also typically easy to categorize, with websites that contain toxic green backgrounds accompanied by blurry photos of George Soros  [^11] https://web.archive.org/web/20240112053008/http://www.endgamethemovie.com/, proclamations of white nationalism  [^12] https://web.archive.org/web/20240108044820/https://www.stormfront.org/forum/, or full links to alien siting websites and paid psychic services  [^13] https://web.archive.org/web/20240116140224/https://rense.com/. Though our webgraph model is content-agnostic, we hypothesize that the webgraph and site attributes will exhibit similarly "easy" patterns across the extremes. To test this hypothesis, we implement a curriculum learning batching procedure Bengio et al. (2009).
Appendix C Partial F1 threshhold sensitivity
@@
To investigate the partial F1 metric we experiment with various classifier confidence thresholds. Naturally, as confidence increases, we see higher precision and lower recall (Figure 5). The disparity between partial recall and partial precision implies that our discovery process is not very precise. However, in the manual evaluation of Precision@k, we find precision is actually very high. This indicates a fundamental issue with the partial precision metric; it underestimates model performance, particularly as the MBFC reference list ages. This is in line with previous findings that unreliable domain lists quickly become outdated (Carragher et al., 2024), and it complicates the evaluation of discovery processes. Further work on domain reliability discovery metrics is needed.
@@
For Precision@k assessments, two annotators were asked “is this website unreliable” and asked to provide a rationale for why or why not. Annotations involved exploration of the domain and lateral reading—looking at what other reliable sources have written about the domain. We publicly release these annotations alongside annotator rationales. [^14] Following acceptance.

### v2

@@ Acknowledgements
Appendix A Appendix A: Additional Data Details
To our knowledge, the list created in (Lin et al. 2023) is the most comprehensive public domain reliability rating list as of August 2024. The domains in this list are ranked by the calculated first principal component where low scores correspond to high expert agreement on unreliability and high scores correspond to expert agreement on reliability. While these scores are very sensible across wide principal component score gaps, locally the relative orderings are less clear. Treating website reliability as continuous also muddies the interpretability and the discovery process, as there’s not a transparent reason a website should have a score of 0.571 as opposed to 0.570. To address this issue, we elected to binarize the website reliability labels. After manual inspection of the data, we elected to consider the bottom two quintiles as unreliable (corresponding to a principal component threshold of 0.5162).
From Ahrefs, we were unable to pull backlinks for 193 of the domains. A random sample of 10 of 193 domains and found 9 of them were either dead or unreachable. We therefore elected to drop these 193 domains, leaving us with 11,327 unique reliability-labeled domains and 32,431 unique unlabeled backlinking domains for a total of 43,758 domains. For each of these 43,758 domains, we pull 23 attributes, which for an individual domain, contains fields like the total number of backlinks and outlinks, number of backlinks coming from .edu or .gov domains, and number of referring pages  [^17] https://ahrefs.com/api/documentation/metrics.
Similar to the backlink network constructed in (Carragher, Williams, and Carley 2024a), there is clear assortativity across label-reliability groups as can be seen in Figure 4. The prevalence of unlabeled nodes in our network overwhelms a node attribute assortativity coefficient calculation $r_{A}=-0.336$. We re-calculate node attribute assortativity on an induced subgraph that only maintains edges between labeled nodes. The subgraph contains less structure and 176 components, but nevertheless displays strong positive associativity $r_{A}=0.376$. In other words, domains of the same (binary) reliability are more likely to link to one another than linking to a domain of a different reliability. This provides a strong justification for the use of network-based models.
Twitter Cleaning
Firstly, we drop all tweets that do not explicitly include a link to the URL. Second, users with fewer than 10 observed tweets over the time period were dropped. To further condense the dataset, we only include tweets which appeared at least 3 times in the dataset. As reposting and retweeting are important influence metrics on Twitter, this drops the tweets that likely did not receive as much attention. We also dropped domains that were only linked to by a single user and users that only linked to a single domain, as we hypothesized these pendulum nodes would be of limited use given the size of the graph. As a result of the expense of the webgraph attribute API, we chose to further restrict the Twitter data to only include tweets that mention at least one of the 43,758 domains for which we extracted attributes. This corresponds to keeping tweets that mention at least one of the 11,327 labeled domains for which we have attributes or tweets that mention one of the 11,327 and co-mention any of the 32,431 unlabeled domains.
Dredge EDA
@@ Appendix A Appendix A: Additional Data Details
We applied the python library langdetect (Danilák 2014) to the list of 3,933 dredge words used in the paper, but we observe that automated language extraction of dredge words is a challenge. We observed many dredge words are disembodied fragments, names of people, or simply foreign words adopted by various communities. For example, several Hindi-language-origin yoga poses appear in the dredge word list: parivrtta anjaneyasana and parivrtta utkatasana. These terms surface gaia, a conspiracy pseudoscience website [^18] https://mediabiasfactcheck.com/gaia/. There are additionally arabic and chinese dredge words that surface creation, a website that frequently promotes pseudoscience [^19] https://mediabiasfactcheck.com/christian-ministries-international/. Qualitatively, hatespeech is fairly uncommon in the dredge words we identified; 11 unique dredge word phrases contain the non-euphemistic writing of the n-word and 3 others contain a homophobic slur.
Appendix B Appendix B: Curriculum EDA
@@
Among the websites rated most reliable, the style, content, and widespread name recognition would very likely result in most annotators to correctly identifying the sites as reliable (e.g., reuters.com, nasa.gov, smithsonianmag.com, nature.com). The landing pages of some of the least reliable labeled domains are also typically easy to categorize, with websites that contain toxic green backgrounds accompanied by blurry photos of George Soros  [^20] https://web.archive.org/web/20240112053008/http://www.endgamethemovie.com/, proclamations of white nationalism  [^21] https://web.archive.org/web/20240108044820/https://www.stormfront.org/forum/, or full links to alien siting websites and paid psychic services  [^22] https://web.archive.org/web/20240116140224/https://rense.com/. Though our webgraph model is content-agnostic, we hypothesize that the webgraph and site attributes will exhibit similarly ”easy” patterns across the extremes. To test this hypothesis, we implement a curriculum learning batching procedure (Bengio et al. 2009).
Appendix C Appendix C: Partial F1 threshhold sensitivity
@@
To investigate the partial F1 metric we experiment with various classifier confidence thresholds. Naturally, as confidence increases, we see higher precision and lower recall (Figure 5). The disparity between partial recall and partial precision implies that our discovery process is not very precise. However, in the manual evaluation of Precision@k, we find precision is actually very high. This indicates a fundamental issue with the partial precision metric; it underestimates model performance, particularly as the MBFC reference list ages. This is in line with previous findings that unreliable domain lists quickly become outdated (Carragher, Williams, and Carley 2024a), and it complicates the evaluation of discovery processes. Further work on domain reliability discovery metrics is needed.
@@
For Precision@k assessments, two annotators were asked “is this website unreliable” and asked to provide a rationale for why or why not. Annotations involved exploration of the domain and lateral reading—looking at what other reliable sources have written about the domain. We publicly release these annotations alongside annotator rationales. [^23] Following anonymity period.